find the polynomila for the perimeter and for the area, if one side is b+4 and the other is b, what is the area and perimeter. Can you help me please? thank you
Answer by CubeyThePenguin(3113)

area = (b + 4)(b) = b^2 + 4b

perimeter = 2(b + 4) + 2b = 4b + 8